The physical fitness of an athlete is often measured by how much oxygen the athlete takes in (which is recorded in milliliters per kilogram, ml/kg). The mean maximum oxygen uptake for elite athletes has been found to be 65 with a standard deviation of 6.1. Assume that the distribution is approximately normal.

(a) What is the probability that an elite athlete has a maximum oxygen uptake of at least 65 ml/kg?

(b) What is the probability that an elite athlete has a maximum oxygen uptake of 60 ml/kg or lower?

(c) Consider someone with a maximum oxygen uptake of 27 ml/kg. Is it likely that this person is an elite athlete? Write "YES" or "NO."
Answer #1

a)
z = (x - μ)/σ
z = (65 - 65)/6.1 = 0

Therefore,
P(X >= 65) = P(z <= (65 - 65)/6.1)
= P(z >= 0)
= 1 - 0.5 = 0.5

b)
z = (x - μ)/σ
z = (60 - 65)/6.1 = -0.82

Therefore,
P(X <= 60) = P(z <= (60 - 65)/6.1)
= P(z <= -0.82)
= 0.2061

c)
z = (x - μ)/σ
z = (27 - 65)/6.1 = -6.23

Therefore,
P(X <= 27) = P(z <= (27 - 65)/6.1)
= P(z <= -6.23)
= 0

As the probability is significantly low, it is not an elite athlete